OpenStudy (anonymous):

Need ans to this question on Wave maths,so that i can compare your solutions to mine. Transverse waves on a string have a speed of 8.10 m/s, amplitude 7.10 cm and wavelength of 0.320 m.D waves travel in neg x-direction, and at t = 0 the x=0 end of the string has its maximum upward displacement. a)Find the frequency, period and wave number of these waves. b)Give a wave function describing these waves. c)Find the transverse displacement of a particle at x=0.35m at time t=0.15s. d)How much time must elapse from the instant in part(c)until d particle at x=0.35m has its next max upward displacementNssemes on a string have a speed of 8.10 m/s, amplitude 7.10 cm and wavelength of 0.320 m.D waves travel in neg x-direction, and at t = 0 the x=0 end of the string has its maximum upward displacement. a)Find the frequency, period and wave number of these waves. b)Give a wave function describing these waves. c)Find the transverse displacement of a particle at x=0.35m at time t=0.15s. d)How much time must elapse from the instant in part(c)until d particle at x=0.35m has its next max upward displacement

OpenStudy (anonymous):

freequency=25.31 hz (f=vel/wavelength ) time period=0.04s (1/f) wave no=1/lambda =3.125 w=2*pi*25.31 k=pi/0.16 so, -0.071sin(kx-wt) c) replace values of x and t d) later...